OpenStudy (anonymous):

Will give Fan and Medal!! The steps to derive the quadratic formula are shown below:

OpenStudy (anonymous):

The question is in file

OpenStudy (anonymous):

@zepdrix @iambatman

zepdrix (zepdrix):

hey c: Hmm I don't see the mistake -_- lemme keep looking..

OpenStudy (anonymous):

I know for sure that the answer is not B, so atleast that narrows it down a bit. And thank you for helping me c:

zepdrix (zepdrix):

It's definitely not A, we have to add that squared term to each side.

zepdrix (zepdrix):

Check out steps 6 and 7, which one do you think it is? c: Where is the error?

OpenStudy (anonymous):

Okay so either C or D

OpenStudy (anonymous):

I beleive that the mistake would be in the last step i think

zepdrix (zepdrix):

Yessss! \[\Large\rm \text{In step 6:}\quad \sqrt{-\frac{c}{a}+\left(\color{orangered}{\frac{b}{2a}}\right)^2}\]\[\Large\rm \text{In step 7:}\quad \sqrt{-\frac{c}{a}+\left(\color{orangered}{\frac{b}{4a}}\right)^2}\]Do you see the boo boo they made?

OpenStudy (anonymous):

Yes, they did 4a instead of 2a, Right?

zepdrix (zepdrix):

Yes. Ok good job \c:/

OpenStudy (anonymous):

Thank you very much c: @zepdrix